OpenStudy (anonymous):

Let A={1,2,3,4,5,6}* {1,2,3,4,5,6} and define a relation R on A as follows: For (a,b),(c,d)ϵA, (a,b)R(c,d) if and only if ab=cd. (i) Verify that R is an equivalence relation on A. (ii) Determine the equivalence class [(2,3)]. Help me pls~ i really duno how to do it~

OpenStudy (anonymous):

it is \(A=\{1,2,3,4,5,6\}\times \{1,2,3,4,5,6\}\)

OpenStudy (anonymous):

ya~ is A={1,2,3,4,5,6}×{1,2,3,4,5,6}

OpenStudy (anonymous):

is the relation given by \((a,b)R(c,d)\iff ad=bc\)

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

not what you wrote, right? what i wrote

OpenStudy (anonymous):

\(ad=bc\) defines the relation, right?

OpenStudy (anonymous):

no~ is ab=cd~

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

for example \((2,6)R(4,3)\)

OpenStudy (anonymous):

because \(2\times 6=4\times 3\)

OpenStudy (anonymous):

so you have to show that this relation is Reflexive, Symmetric Transitive do you know what these mean and how to do it?

OpenStudy (anonymous):

i am just wondering. if you do not, that if fine i can try to explain

OpenStudy (anonymous):

ok~ the reflexive, symmetric and transitive i know how to do it~ how about the (ii)? bcz i don't the question asking what.

OpenStudy (anonymous):

that is the easy part!!

OpenStudy (anonymous):

if you have an equivalence relation \(\equiv\) then it is like having an equal sign what you have to do is find all the pairs that are in the equivalence class of \((2,3)\) that is all pairs \((c,d)\) such that \(c\times d=2\times 3=6\)

OpenStudy (anonymous):

for example \((6,1)\) is in that equivalence class, since \(6\times 1=6\)

OpenStudy (anonymous):

i think there are exactly 4 ordered pairs in that class does this make sense? you did all the hard work if you proved the rst part

OpenStudy (anonymous):

oh~ i get what you mean~ okok~ thx~~ =D

OpenStudy (anonymous):

yw good luck for further understanding, find all the equivalence classes. it is not hard